OpenStudy (anonymous):

After 85% reduction you purchase a new washing machine for $75. What was the original price of the machine

OpenStudy (anonymous):

.15X=75 Solve for X .15 cause that is what % you are paying. X is original price. 75 is product.

OpenStudy (anonymous):

x=500?

OpenStudy (anonymous):

yes is 500

OpenStudy (kropot72):

Let x = original price of machine x - 0.85x = 75 x(1 - 0.85) = 75 0.15 * x = 75 \[x=\frac{75}{0.15} = 500\] The original price was $500